# Operationalizing content moderation “accuracy” in the Digital Services Act

### v3

@@ 3.3. Metrics and their legal implications
Having opened up “accuracy” to the possibility of broader interpretation, we consider its operationalization with basic accuracy metrics from computer science, and discuss their legal implications.
@@
This metric is by far mentioned the most in EU law, and technically defined as: $\text{Accuracy}=\frac{TP+TN}{TP+TN+FP+FN}$ where TP, FP, TN, FN are true positives, false positives, true negatives, false negatives, respectively. Using accuracy as a metric for content moderation is problematic because of label-imbalance in the content moderation setting: in a huge set of content like social media posts, the illegal posts are usually far outnumbered by the posts that do not violate specific laws. Accuracy can be highly inflated if the test data is label-imbalanced, as a classifier that only outputs the majority label will achieve high accuracy. In §5.1, we show that even a simple classifier can achieve an accuracy of 95% due to label imbalance in the data set. Alternatively, precision and recall are separate accuracies of the predicted positives and true positives, respectively, which remain informative in label imbalanced settings.
@@
(1) $\text{Precision}=\frac{TP}{TP+FP}.$
As each FP is an individual violation to the freedom of expression (by failing to allow protected free speech), precision is a probabilistic reflection of the moderation system’s respect to the right of freedom of expression. A low precision would indicate that much content which is not actually violating is being positively identified and acted against. If a social media platform exceeds their legal obligation of content moderation, their overmoderation interferes with the right to freedom of expression. It would only be justified if it does not present a disproportionate interference to this right, which is measured by precision.
@@
(2) $\text{Recall}=\frac{TP}{TP+FN}.$
@@
The DSA has to balance the freedom of expression with the right to non-discrimination, and such a balance would be reflected mathematically by the precision-recall tradeoff. [^16] https://scikit-learn.org/stable/auto_examples/model_selection/plot_precision_recall.html A probabilistic classifier (i.e. one that predicts the probability of whether a given piece of content is illegal, and takes an action against it, such as removal, if the score is greater than a decision threshold) can trade off precision for recall and vica versa by tuning its decision threshold. By increasing the threshold, the classifier is more conservative in its positive predictions, which increases the precision at the cost of recall. Likewise, precision can be traded off for recall by decreasing the threshold. The decision threshold is an implicit but intentional setting set by the platforms that deploy these classifiers (Douek, 2021b). Regulators should recognize that, while improving both precision and recall simultaneously is non-trivial, trading them off is a technical option to strike the appropriate balance between fundamental rights.
TP, FP, TN, FN are atomic
There are a number of metrics which can be computed when considering content moderation as a classification problem (which is the perspective the DSA takes). Metrics such as F1, which is the harmonic mean of the precision and recall, can serve as useful summary statistics. However, requiring only summary statistics will obscure whether platforms are over- or undermoderating. As FN and FP directly correspond to the interference of Fundamental Rights (listed in §3.1), we encourage regulators to reason about precision and recall, which are their normalized measures. It is sufficient to require the reporting of TP, FP, TN, and FN, from which precision, recall, F1, and other metrics can be calculated.
4. Statistically estimating legal reporting requirements
Given that interpreting content moderation “accuracy” as precision and recall is legally sensible, we now discuss the technical implementation of precision and recall. Computing such metrics can be non-trivial, as we cannot exhaustively annotate the subsets (i.e. all visible content) needed to compute the desired quantities. The only practical alternative is statistical estimation of these metrics via sampling and annotation. We highlight two statistical concepts necessary for the legal purposes of transparency reporting:
Unbiasedness.
For legal purposes, any estimate we provide should be “true” and reflect the true value of the metric. In statistical terms, we can think of the true value of the metric as one computed over a set of exhaustive human annotations (the “expectation”, see eq. 3). Since exhaustive annotation may not be possible, we sample and annotate to compute an estimate. A statistical estimate comes in the form of a point estimate (best guess of true value) and a confidence interval (i.e. we are 95% confident that the true value is between this range) and the estimator is unbiased if the point estimate, averaged over many experiments, is equal to the true value. To achieve this property requires human ground truth annotations, as solely relying on classifiers introduces bias because they make errors relative to the ground truth (Gorwa et al., 2020).
Number of samples (cost).
An unduly burdensome reporting requirement would interfere disproportionately with the right to conduct business (Art. 16 CFR). Thus, our statistical analyses focuses on cost: how many annotations are needed to provide a reasonable width confidence interval? In particular, recall estimation is costly because it requires estimating the number of remaining positives among all visible content (false negatives, $FN$). The rarer they are, the more samples are needed to construct a reasonable estimate for the false negatives. This same challenge is well known for tasks like information retrieval, where recall is defined over all webpages on the web (Aslam et al., 2006). With variance reduction techniques, we can provide a narrower interval for a fixed cost (Owen, 2013).
4.1. Estimating precision
@@
Auditing content moderation precision is qualitatively difficult but statistically simple, and so is not the focus of our work. In short, precision can be estimated well by annotating a random sample of detected content. Appendix B contains a discussion on estimating precision.
4.2. Estimating recall
@@
Measuring recall poses a challenging statistical problem, which is the focus of our work. Unlike precision, recall requires looking through the predicted negatives (i.e. visible content on the platform during the last reporting period) to determine the number of false negatives $FN$. Typically, the prevalence of false negatives among the negatives is rare e.g. $0.10<p<<$ 0.1. Appendix A.1 contains a statistical formulation estimating the prevalence $p$. The simplest way to estimate this probability is by applying the random sampling estimator $\hat{p}$, where a random sample of negatives is drawn and annotated to estimate the prevalence of false negative. When estimating the probability of a rare event, we would like to control the coefficient of variance $\textsc{CV}_{\hat{p}}=\textsc{SE}_{\hat{p}}/\hat{p}$ which is the relative estimation error to the prevalence. In other words, we are estimating the rare probability within some percentage accuracy (Dixon et al., 2005). If we require estimates to have a $\textsc{CV}_{\hat{p}}\leq 20\%$, the rarer the true prevalence, the smaller the standard error will need to be.
Propagating confidence intervals
Once we have a statistical estimate for the false negative rate $p$, the confidence intervals on $p$ can propagate to a confidence interval on recall. In the case where we know the number of true positives $TP$, the confidence intervals on $p$ can directly be translated to confidence intervals on $FN$ then to recall, by plugging in the upper and lower estimates of $p$ into the recall equation. If $TP$ is not a known quantity and also an estimate, confidence intervals can be obtained through bootstrap simulation.
4.3. Stratified sampling
To reduce the burden of a recall reporting requirement, we propose applying stratified sampling, which is a variance reduction technique. Stratified sampling can reduce the cost of recall estimation, by combining it with a learned classifier for stratification. Crucially, it is unbiased and more efficient than a random sampling estimator. Stratified sampling is unbiased even if the classifier which partitions the strata has undesirable properties such as low accuracy or low fairness (Borkan et al., 2019), as a human is the final decision maker. A suboptimal classifier here may cause the method to require more annotations, but it cannot introduce bias. Appendix A.2 contains notation for stratified sampling.
Intuition.
Given any partition of the data into strata (i.e., bins), we can apply the stratified sampling estimator. This estimator is computed by annotating a random sample within each stratum, then combining the mean for each stratum in an unbiased manner. A good stratification can significantly improve the variance of the stratified sampling estimator over random sampling. In particular, if some strata have very low or very high prevalance, those strata will have lower variance. Then, we can allocate our samples to focus on estimating prevalences of the other high variance strata.
4.4. Binning and allocation
@@
The variance of stratified sampling depends on two choices: binning (i.e., how to partition the examples) and allocation (i.e., how many examples to annotate in each stratum). We now describe methods to choose each.
@@
An effective way to create strata is with a binning classifier trained to predict the probability that a given piece of content violates community guidelines. With the predictions scores from the binning classifier, content can be binned in a few ways by their score. The number of bins $L$ is a hyperparameter, which we investigate in §5. (1) Equal width binning. The binning classifiers we consider score each comment from $[0,1]$. Equal-width binning splits this interval into $L$ equal width intervals each corresponding to a bin. Sampling across bins thus samples across the entire scoring range of the classifier. (2) Quantile binning. Once the comments are scored, quantile binning splits the scoring range into equal sized quantiles, where each bin will contain the same number of points. The range of scores within each quantile will depend on the scoring distribution of the classifier. (3) Oracle binning. This binning is not possible in practice as it requires the labels of all the data in the pool, but we include it as an oracle method. Using a recursive search procedure and the labels of all the training data, we can search for a good binning. At each recursive step, a bin is split into two at a point where the resulting variance is minimized. We recursively break down the binning classifier’s scoring interval into a number of bins that is a power of two.
@@
4.5. Analyzing the burden of estimation
The variance of the random and stratified sampling estimators can be calculated analytically, so their power (i.e. number of samples needed to achieve some estimation accuracy) can also be calculated if we assume the central limit theorem. This assumption is justified because the sample sizes we consider are large, so the estimators are normally distributed.
We consider the goal of reporting $\hat{p}\pm(0.1)\hat{p}$ with 95% probability. The coefficient of variance is effectively required to be $\textsc{CV}_{req}=0.1/z_{.95}$, and in turn $\textsc{SE}_{req}=(0.1)\hat{p}/z_{.95}$, where $z_{(1-\alpha)}$ is test statistic of the normal variable. Since $\textsc{SE}_{req}$ is a function of $n$, the total number of annotations collected, we can solve for $n$ for each estimator. Power calculations for random sampling are given in Appendix A.4. Appendix A.4 and A.5 contain closed form solutions for the cost of random and stratified sampling, respectively.
5. Efficiently estimating recall for a keyword filter
In this section, we determine which choices of binning, allocation (described in §4.4), and classifier are best for estimating recall with stratified sampling. We use CivilComments (Borkan et al., 2019) as our testbed and emulate a content moderation system by constructing a keyword-based filter. While Civilcomments is annotated for toxicity and it is not an operationalized legal definition or community guideline, this setting provides a large dataset to study recall estimation in general. Similar to pool-based active learning (Settles, 2012), we assume in our simulated experiments that the dataset is initially unlabeled, and we “annotate” an example by revealing the label from the dataset.
5.1. A simple keyword filter
@@
In total, the dataset contains about 100K toxic comments out of 1.7M comment, and the overall prevalence of toxicity is 5.9%. The filter achieves a precision of 58%, removing 57K comments with 32K of these actually being toxic. Amongst the unfiltered examples, the prevalence is 4.1%. The filter would only achieve a recall of 33%, removing about 33K of the total 100K toxic comments. However, the accuracy would be nearly 95%.
@@
Figure 1. Number of samples required to estimate the prevalence of toxicity (4.1%) within 20% in the unfiltered CivilComments, with different number of strata. Binning is based on predicted scores from a finetuned Roberta model. All pilot results are averaged over 30 trials. As the number of strata grows, estimators using optimal allocation require fewer samples. Pilot methods outperform the equal allocation baseline, but when there are too many strata, annotation effort is wasted on the pilot samples.
5.2. Experimental setup
@@
The CivilComments dataset (Borkan et al., 2019) contains comments labeled by their toxicity, from the archive of the Civil Comments platform, a commenting plugin for independent news sites. Each comment is annotated for toxicity by at least 10 annotators, and if half of the annotators label it toxic, the toxicity label is positive. For our study, we use 100K random examples of the training set for training, and the rest of the 1.7M training examples for simulation. This creates a large testing setting which better represents content moderation, where $\mathcal{U}$ may be large and the number of examples to train the model is small. This dataset naturally presents a rare prevalence estimation problem as $5.9\%$ of comments are toxic.
@@
Roberta (Liu et al., 2019) is a transformer-based language model which is pretrained on a large corpus of English data, which we find to have strong performance for CivilComments after finetuning. We use typical hyperparameters for small datasets (Mosbach et al., 2021), and make two critical adjustments: increasing the batch size to 32 (we hypothesize that since the labels can be noisy, the gradients are unstable), and subsampling the negative examples so that the training data is balanced (the resulting training data has 11k examples).
5.3. Discussion

### v5

@@ 3.3 Metrics and their legal implications
Having opened up “accuracy” to the possibility of broader interpretation, we consider its operationalization with basic accuracy metrics, and discuss their legal implications.
TP, FP, TN, and FN
TP, FP, TN, and FN are the true positives, false positives, true negatives, and false negatives, respectively. This section discusses content moderation in terms of classifying illegal content (which is a perspective the DSA takes, see §3.1). In this context, TP and TN represent correctly identified illegal and free speech. Each FP is an individual violation to the freedom of expression (by misclassifying and acting against free speech), while each FN is a failure to identify and act against discriminatory content.
TP, FP, TN, and FN are primitives from which classification metrics are derived. Therefore, the reporting of TP, FP, TN, and FN, would be sufficient to calculate accuracy, precision, recall, F1, or any other metric. As FN and FP directly correspond to the interference of Fundamental Rights (listed in §3.1), we encourage regulators to reason about precision and recall, which are their normalized measures. Metrics such as accuracy or F1 (the harmonic mean of the precision and recall) can serve as useful summary statistics. However, only reporting summary statistics will obscure whether platforms are over- or undermoderating.
@@
This metric is by far mentioned the most in EU law, and defined as: $\text{Accuracy}=\frac{TP+TN}{TP+TN+FP+FN}$which is the number of correct predictions divided by the total. Using accuracy as a metric for content moderation is problematic because of label-imbalance in the content moderation setting: in a huge set of content like social media posts, the illegal posts are usually far outnumbered by the posts that do not violate specific laws. Accuracy can be highly inflated if the test data is label-imbalanced, as a classifier that only outputs the majority label will achieve high accuracy. In §5.1, we show that even a simple classifier can achieve an accuracy of 95% due to label imbalance in the data set. Alternatively, precision and recall are separate accuracies of the predicted positives and true positives, respectively, which remain informative in label imbalanced settings.
@@
$\text{Precision}=\frac{TP}{TP+FP}.$ (1)
As each FP is an individual violation to the freedom of expression (by misclassifying and acting against free speech), precision is a probabilistic reflection of the moderation system’s respect to the right of freedom of expression. A low precision would indicate that much content which is not actually illegal is being positively identified and acted against. If a social media platform exceeds their legal obligation of content moderation, overmoderation would interfere with the right to freedom of expression. Content moderation is only justified if it does not present a disproportionate interference to this right, which is measured by precision.
@@
$\text{Recall}=\frac{TP}{TP+FN}.$ (2)
@@
The DSA has to balance the freedom of expression with the right to non-discrimination, and such a balance would be reflected mathematically by the precision-recall tradeoff. [^18] https://scikit-learn.org/stable/auto˙examples/model˙selection/plot˙precision˙recall.html A probabilistic classifier (i.e. one that predicts the probability of whether a given piece of content is illegal, and takes an action against it, such as removal, if the score is greater than a decision threshold) can trade off precision for recall and vica versa by tuning its decision threshold. By increasing the threshold, the classifier is more conservative in its positive predictions, which increases the precision at the cost of recall. Likewise, precision can be traded off for recall by decreasing the threshold. The decision threshold is an implicit but intentional setting set by the platforms that deploy these classifiers, and requiring the reporting of precision and recall would make clear the effects of this implicit threshold (Douek 2021b). Regulators should recognize that, while improving both precision and recall simultaneously is non-trivial, trading them off is a technical option to strike the appropriate balance between fundamental rights.
4 Statistically estimating legal requirements
Given that interpreting content moderation “accuracy” as precision and recall is legally sensible, we now discuss the technical implementation of precision and recall. This section starts with an example of statistical estimation intended for an interdisciplinary audience.
As an example, suppose a social media platform deploys an automated content moderation system that deletes content which violates their community guidelines. We may want to estimate the prevalence (or proportion) of content violations which remain visible on the platform. This prevalence would be the false negative rate of the deployed moderation system, and can be used to calculate the recall (more details in §4.2). We outline three components of statistical estimation here:
Expectation
Statistical estimation begins with positing the ideal quantity we are estimating. Ideally, an annotator could label every single piece of visible content on the platform to determine whether it is violating or not. We could then directly calculate the false negative rate, which we denote as $p$ (where $0\leq p\leq 1$). Since $p$ is derived from an exhaustive average, $p$ is called an expectation in statistical terms. However, social media platforms accumulate millions of pieces of content a day, so exhaustive annotation would not be feasible. Therefore, the true $p$ will be unknown.
Sampling
The only practical alternative is to estimate the false negative rate $p$ by sampling some visible content and then annotating this sample. The prevalence of content in the sample, which we denote as $\hat{p}$, can be used as an estimate for $p$. The number of samples directly relates to the implementation cost: each sample needs to be labelled by an annotator.
Standard error
A better estimate $\hat{p}$ is one that is close to $p$, on average. The average distance between $\hat{p}$ and $p$ is a standard error in statistical terms. If the estimate $\hat{p}$ is unbiased, more samples results in smaller standard error. Estimation error is then a function of these two components:
  • Unbiasedness For legal reporting, we should provide a statistically unbiased estimate $\hat{p}$ which, on average, reflects the false negative rate $p$. When estimating $\hat{p}$, two details ensure statistical unbiasedness. The first is using a “representative” sampling method like random sampling. Stratified sampling is also unbiased by emulating a “representative” sample (see §4.3). Second, human ground truth annotations are required. Solely relying on classifiers introduces bias because they make errors relative to the ground truth (Gorwa, Binns, and Katzenbach 2020).
  • Number of samples (cost) An unduly burdensome reporting requirement would interfere disproportionately with the right to conduct business (Art. 16 CFR). Thus, our statistical analyses focuses on cost: how many annotations are needed to provide a reasonably good estimate? In general, a larger sample yields better estimates. On the other hand, as $p$ approaches 0, it requires more samples to accurately estimate e.g. if $p=0.02$, 10 samples would likely yield estimates $\hat{p}=0.0$ or $\hat{p}=0.1$, where the sample contains 0 or 1 pieces of violating content. More samples would be needed to refine the granularity of $\hat{p}$.
4.1 Estimating precision
@@
Auditing content moderation precision is qualitatively difficult but statistically simple, and so is not the focus of our work. In short, precision can be well estimated by annotating a random sample of positively predicted content. Appendix B contains a discussion on estimating precision.
4.2 Estimating recall
Measuring recall poses a challenging statistical problem, which is the focus of our work. To compute recall, both TP and FN are needed. TP can be computed from the precision (the positives multiplied by the precision), but estimating FN requires looking through the predicted negatives of the content moderation system (i.e. visible content on the platform during the last reporting period) to determine the number of false negatives. Similarly, recall is challenging to measure for tasks like information retrieval, as the false negatives are the relevant webpages missed by the retrieval system across the web (Aslam, Pavlu, and Yilmaz 2006).
@@ 4.2 Estimating recall
Typically, the prevalence of misclassified violations amongst all content is rare, e.g. 1 in 1000, so estimating the false negative rate requires a large number of samples to provide a reasonable estimate. Appendix A.1 contains the statistical formulation on estimating the false negative rate $p$. The simplest way to estimate $p$ is to apply the random sampling estimator, where a random sample of negatives is collected and annotated to compute $\hat{p}$ as an estimate for $p$. When estimating a small value like $p$, we want the estimation error to be small, relative to $p$. In other words, we are estimating the rare probability $p$ within some percentage accuracy. In statistical terms, we want to ensure the coefficient of variance $\textsc{CV}_{\hat{p}}=\textsc{SE}_{\hat{p}}/\hat{p}$ is low i.e. $\textsc{CV}_{\hat{p}}\leq 20\%$, where $\textsc{SE}_{\hat{p}}$ is the standard error of estimate $\hat{p}$ (Dixon, Ellison, and Gotelli 2005).
4.3 Stratified sampling
To reduce the burden of recall reporting, we propose using stratified sampling instead of random sampling. Stratified sampling can provide a better estimate with the same number of samples (Owen 2013). Crucially, it is unbiased and more efficient than a random sampling estimator. Appendix A.2 contains notation for stratified sampling and only a high level discussion is given here. The procedure is as follows: a learned classifier is first used to partition e.g. the visible comments into bins. Then, the estimator is computed by annotating a random sample within each stratum, and combining the prevalences across stratum in an unbiased manner.
Intuition
A good stratification can significantly increase the efficiency of the stratified sampling estimator over random sampling. If some strata have a very low or very high prevalence of violating comments, few samples are needed in those strata to estimate the prevalences well. This allows us to allocate our annotation effort on estimating strata with medium prevalences (which have higher variance). Stratified sampling is unbiased even if the stratification classifier has undesirable properties such as low accuracy or low fairness (Borkan et al. 2019), as a human is the final decision maker. A suboptimal classifier may cause the estimator to be less efficient, but it cannot introduce statistical bias.
4.4 Binning and allocation
@@
The variance of stratified sampling depends on two choices: binning (i.e., how to partition the examples) and allocation (i.e., how many examples to annotate in each stratum). We now provide technical details for each choice.
@@
An effective way to create strata is with a binning classifier trained to predict the probability that a given piece of content violates community guidelines. With the predictions scores from the binning classifier, content can be binned in a few ways by their score. The binning classifiers gives each comment a score from $[0,1]$. The number of bins $L$ is a hyperparameter, which we investigate in §5. (1) Equal width binning. Equal-width binning splits this interval into $L$ equal width intervals each corresponding to a bin. Sampling across bins thus samples across the entire scoring range of the classifier. (2) Quantile binning. Once the comments are scored, quantile binning splits the scoring range into equal sized quantiles, where each bin will contain the same number of points. The range of scores within each quantile will depend on the scoring distribution of the classifier. (3) Oracle binning. This binning is not possible in practice as it requires the labels of all the data in the pool, but we include it as an oracle method. Using a recursive search procedure and the labels of all the training data, we can search for a good binning. At each recursive step, a bin is split into two at a point where the resulting variance is minimized. We recursively break down the binning classifier’s scoring interval into a number of bins that is a power of two.
@@
4.5 Analyzing the burden of estimation
The statistical analyses here are focused on cost i.e. how many annotations are needed to provide a good estimate? The number of annotations needed to achieve a fixed estimation error can be calculated analytically for both random and stratified sampling estimators. A few technical details are provided here and we defer the rest to the appendix. Appendix A.4 and A.5 contain closed form solutions for the cost of random and stratified sampling, respectively. To give some intuition, the cost of estimating different prevalences with random sampling are given in Appendix A.4.
We consider the goal to report $\hat{p}$ to within $20\%$ of $p$, with 95% probability. In statistical terms, we calculate the number of samples required for the 95% confidence interval of $\hat{p}$ to be a fixed width relative to $p$ [^19] Once we have a statistical estimate $\hat{p}$ for the false negative rate $p$, the confidence intervals on $\hat{p}$ can propagate to a confidence interval on recall. In the case where we know the exact number of true positives $TP$, the confidence intervals on $p$ can directly be translated to confidence intervals on $FN$, and then to recall, by plugging in the upper and lower estimates of $p$ into the recall equation. If $TP$ is not a known quantity and also an estimate, confidence intervals can be obtained through bootstrap simulation. The coefficient of variance is effectively required to be $\textsc{CV}_{req}=0.2/z_{.95}$, and in turn the standard error is required to be $\textsc{SE}_{req}=(0.2)\hat{p}/z_{.95}$, where $z_{(1-\alpha)}$ is test statistic of the normal variable. Since $\textsc{SE}_{req}$ is a function of the total number of annotations collected, we can solve for this number for each estimator.
5 Efficiently estimating recall for a filter
In this section, we determine which choices of binning, allocation (described in §4.4), and classifier are best for estimating recall with stratified sampling. The efficient and robust estimator derived in this section is used to estimate a content moderation recall for Reddit in §6, where we will continue to provide legal analysis on areas of underspecification in the “accuracy” reporting requirement.
We use CivilComments (Borkan et al. 2019) as our testbed and emulate a content moderation system by constructing a keyword-based comment filter. While Civilcomments is annotated for toxicity and is not an operationalized legal definition or community guideline, this setting provides a large dataset to study recall estimation on a related task. Similar to pool-based active learning (Settles 2012), we assume in our simulated experiments that the dataset is initially unlabeled, and we “annotate” whether an example is a false negative by revealing the label from the dataset.
5.1 A simple keyword filter
@@
In total, the dataset contains about 100K toxic comments out of 1.7M comments, and the overall prevalence of toxicity is 5.9%. The filter achieves a precision of 58%, removing 57K comments with 32K of these actually being toxic. Amongst the unfiltered examples, the prevalence is 4.1%. The filter would only achieve a recall of 33%, removing about 33K of the total 100K toxic comments. However, the accuracy would be nearly 95%.
@@
5.2 Experimental setup
@@
The CivilComments dataset (Borkan et al. 2019) contains comments labeled by their toxicity, from the archive of the Civil Comments platform, a commenting plugin for independent news sites. Each comment is annotated for toxicity by at least 10 annotators, and if half of the annotators label it toxic, the toxicity label is positive. For our study, we use 100K random examples of the training set for training, and the rest of the 1.7M training examples for simulation. This creates a large testing setting which better represents content moderation, where the number of examples to train the model is small but it is used for a large number of predictions. This dataset naturally presents a rare prevalence estimation problem as $5.9\%$ of comments are toxic.
@@
Roberta (Liu et al. 2019) is a transformer-based language model pretrained on a large corpus of English data, which has strong performance for CivilComments after finetuning. We use typical hyperparameters for small datasets (Mosbach, Andriushchenko, and Klakow 2021), and make two critical adjustments: increasing the batch size to 32 (we hypothesize that since the labels can be noisy, the gradients are unstable), and subsampling the negative examples so that the training data is balanced (the resulting training data has 11k examples).
5.3 Discussion
Figure 1: Number of samples required to estimate the prevalence of toxicity (4.1%) within 20% in the unfiltered CivilComments, with different number of strata. Binning is based on predicted scores from a finetuned Roberta model. All pilot results are averaged over 30 trials. Dotted line represents random sampling. As the number of strata grows, estimators using optimal allocation require fewer samples. Pilot methods outperform the equal allocation baseline, but when there are too many strata, annotation effort is wasted on the pilot samples.